A sphere of homogenous linear dielectric material is placed in an otherwise uniform elctric fild E not . find the elctric field inside the sphere?
1
Expert's answer
2019-12-19T11:37:04-0500


Element surface


"dS=R^2sin(\theta)d\theta d\alpha"

Electric field

"E_{add}=k\frac{\sigma dS}{R^2}=k\frac{Pcos(\theta)R^2sin(\theta)d\theta d\alpha}{R^2}=kPcos(\theta)sin(\theta)d\theta d\alpha"

"E_{||}=dE_{add}cos(\theta)=kPcos^2(\theta)sin(\theta)d\theta d\alpha"

Integration

"E_{add}=kP\int cos^2(\theta)sin(\theta)d\theta\int d\alpha=\frac{1}{3}\frac{P}{\varepsilon_0}"

Then the electric field in center of the sphere


"\vec{E}=\vec{E_0}+\frac{1}{3}\frac{\vec{P}}{\varepsilon_0}"